# Exploring the Non-Markovian Dynamics in Depolarizing Maps

### v1

@@ III.2 The negativity of the decay rate
In some cases of interest, the open system dynamics may be written in terms of a time local master equation involving time-dependent functions as prefactors with otherwise Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) form [33, 34]. Then, for some times during the system evolution, negative decay rates may show up, indicative of the non-Markovianiaty of the dynamics. The generalised Markovian dynamics appears when the master equation takes the quasi-GKSL-form [35]
@@
written in a canonical form [19], where the $L_{i}(p)$ are trace-less orthonormal operators. The decay rates $\gamma_{i}(p)\geq 0$ defines a CP-divisible dynamical map. It is natural to regard dynamical maps $\Phi(t,0)$ with master equations of type Eq. (III.2) for which $\gamma_{i}(p)<0$ for some $i$ and some time as candidates for non-Markovian quantum dynamics [35]. In these cases, the dynamical map is no longer CP-divisible.
@@
The Kraus representation Eq. (II) is a solution to the master equation describing the depolarizing map in the canonical form:
@@
Figure (3) illustrates a plot of decay rates $\gamma_{X,Y,Z}(p)$ as a function of $p$ for $\alpha=0.0$ (red dashed curve) and $\alpha=0.7$ (blue solid curve). We observe that when the non-Markovian (perturbation) parameter, $\alpha$, is equal to zero, the decay rates become positive for the entire range of $p$ and then the dynamics is CP-divisible, as expected. Furthermore, when $\alpha=0.7$, the decay rates turn into negative at the singularity point, $i.e.,~{}p\approx 0.78$, indicating non-Markovian dynamics.
Using the canonical decay rates, $\gamma_{i}(p)$, and the master equation, Eq. (18), a function $f_{i}(p)$ is defined [19],
$f_{i}(p)=\{-\gamma_{i}(p),0\}.$ (20)
If $f_{i}(p)=0$, at any time, then the evolution is CP-divisible. The functions $f_{i}(p)$ can be used to quantify the non-Markovianity as
@@
where $\gamma(p)$ is as in Eq. (19). It diverges because of the singularity of $\gamma(p)$ at $p\approx 0.8$ (see Fig. (3). One solution to this problem, following an idea proposed in [7, 20], is to replace $\gamma(p)$ by its normalized form
@@
=$=\Biggl{\{}\log\left(2p(-2+\alpha)-4\alpha+3p^{2}\alpha\right)$
$+\frac{6\alpha(tanh^{-1}\left(\frac{-2+\alpha+3p\alpha}{\sqrt{4-4\alpha+13\alpha^{2}}}\right))}{\sqrt{4-4\alpha+13\alpha^{2}}}\Biggr{\}}^{p=1}_{p=\alpha_{-}}.$ (23)
Figure (4) represents the values of $\mathcal{N}^{{}^{\prime}}_{HCLA}$ (blue, solid curve) as a function of non-Markovian (perturbation) parameter $\alpha$. The monotonic increase of this measure with $\alpha$ indicates non-Markovianity. These results are connected to the RHP measure, $\mathcal{N}_{RHP}$, since both coincide for the two level system [19].
Figure 4: (Color online) A plot of $BLP$ measure $\mathcal{N}_{BLP}$ (red, dashed line), and $HCLA$ measure $\mathcal{N}^{{}^{\prime}}_{HCLA}$ (blue, solid curve).
@@ III.3 The distinguishability of quantum states
The BLP method observes the distinguishability between two initial states, $\rho^{S}_{1}$ and $\rho^{S}_{2}$ by calculating the $trace$ $distance$, $D$, between them. $D$ is a metric quantity on the space of physical states, satisfying $0\leq D\leq 1$, where $D=1$ if and only if $\rho^{S}_{1}$ and $\rho^{S}_{2}$ have orthogonal support, and characterizes a realistic upper bound for the distinguishability between the probability distributions resulting from measurements executed on the quantum states. Therefore, the trace distance can be given an interpretation as the distinguishability between two quantum states [5, 3]. If $\Phi(t,0)$ is CP map, then, it is “CP-divisible” and furthermore is a contraction for the trace distance, $i.e.,$ $D$ is non-increasing under CP-divisible maps,
$D\left[\Phi(t,0)(\rho^{1}_{S}),\Phi(t,0)(\rho^{2}_{S})\right]\leq D\left[\rho^{1}_{S},\rho^{2}_{S}\right],$ (24)
where in general,
$D\left[\rho^{2}_{S},\rho^{1}_{S}\right]=\frac{1}{2}||\rho^{2}_{S}-\rho^{1}_{S}||_{1}=\frac{1}{2}Tr\sqrt{(\rho^{2}_{S}-\rho^{1}_{S})^{2}}.$ (25)
The violation of Eq. (24) would be an indicator of non-Markovian behavior [5]. As an example, we calculate the trace distance between two different initial states, with orthogonal support, which evolve under depolarizing dynamics, Eq. (III.1). Let, $\rho^{1}_{S}(0)=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix},$ and $\rho^{2}_{S}(0)=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$. By applying the depolarizing channel, Eq. (III.1), on these initial states, the final states turn out to be
$\rho^{1}_{S}(t)$ $=\Phi(t,0)\left[\rho_{S}^{1}(0)\right]$
$=\frac{1}{2}\begin{pmatrix}\frac{1}{2}&\frac{p}{2}+\frac{\alpha p}{2}-\frac{3}{8}\alpha p^{2}-\frac{1}{2}\\ \frac{p}{2}+\frac{\alpha p}{2}-\frac{3}{8}\alpha p^{2}-\frac{1}{2}&\frac{1}{2}\end{pmatrix},$ (28)
$\rho^{2}_{S}(t)$ $=\Phi(t,0)\left(\rho_{s}^{2}(0)\right)]$
@@
$D\left[\rho^{2}_{S},\rho^{1}_{S}\right]$ $=\frac{1}{2}Tr\sqrt{(\rho^{2}_{S}-\rho^{1}_{S})^{2}}$
@@
Figure 5: (Colour online) Log plot of trace distance $D$ for $\alpha=0.0$ (solid red curve), $\alpha=0.7$ (blue dashed curve), and $\alpha=0.9$ (orange dashed-dotted curve), as a function of $p$. Note that, $\alpha=0.0$ describes the Markovian dynamics since $D$ decreases monotonically for all values of $p$.
The log plot of trace distance $D$ as a function of $p$ with $\alpha=0$ (solid red curve), $\alpha=0.7$ (blue dashed curve), and $\alpha=0.9$ (orange dashed-dotted curve), under the considered non-Markovian depolarizing noise is depicted in Fig. (5). Note that larger $\alpha$ shows a larger enhancement region, suggesting larger non-Markovianity in the sense of BLP [5]. The BLP measure of non-Markovianity, $\mathcal{N}_{BLP}$, is given by
@@
If $\frac{dD}{dp}>0$ for some $p$, then the dynamics is non-Markovian [7], and the quantification of this non-Martkovianity can be done by calculating $\mathcal{N}_{BLP}$
@@
$\mathcal{N}_{BLP}$ is illustrated in Fig. (4) (red, dashed line) and shows an agreement with the quantification of non-Markovianity according to the normalised $\mathcal{N}^{{}^{\prime}}_{HCLA}$.
@@ III.4 Using quantum memory witness
The intermediate dynamical map $\Phi(t,s)$ and the corresponding Choi matrix $\chi(\alpha,q,p)$ are useful for the observation and certification of non-Markovian memory effects. The breaking of CP divisibility can be captured using the Choi matrix, which is Hermitian, allowing one to write a spectral decomposition as $\chi(\alpha,q,p)=\sum_{i}\lambda_{i}P_{i}$. The $P_{i}$ are orthogonal projections derived using corresponding eigenvectors, and the $\lambda_{i}$ are eigenvalues associated with $\chi(\alpha,q,p)$. The occurrence of negative $\lambda_{i}$ points to non-Markovianity. Following this, the Bell states (|$(\ket{\phi^{+}},\ket{\psi^{\pm}})$ are observed to serve as witness operators for depolarizing channel [37] giving
$tr(\ket{\phi^{+}}\bra{\phi^{+}}\chi(\alpha,q,p))=\Lambda_{I},$
$tr(\ket{\psi^{\pm}}\bra{\psi^{\pm}}\chi(\alpha,q,p))=\Lambda_{i=x,y,z}.$ (35)
One of the eigenvalues ($\Lambda_{I},\Lambda_{i=x,y,z}$) becomes negative for $q=0.3$ and $\alpha>0.7$, indicating that the intermediate evolution is not CP divisible in the interval [$q,p$] and hence non-Markovian. The map admits information backflow in this interval, and usual tools lack the ability to distinguish this information as classical or quantum in nature. We use [21] to characterise the quantum nature of information backflow in the intermediate depolarizing map. The existence of the Hermitian witness operator, in the form of the Bell states, then allows the establishment of the variable $X_{\chi(\alpha,q,p)}$ corresponding to $\chi(\alpha,q,p)$. This is used to ascertain the quantum nature of the information backflow, characterized by non-monotinicity of $X_{\chi(\alpha,q,p)}$, by satisfying the following condition [21]
@@
where $\vec{s}$ consists of $s_{i}=tr(\chi(\alpha,q,p)(I\otimes\sigma_{i}))$, and $T$ is a 3$\times$3 matrix consisting of elements $T_{ij}=tr(\chi(\alpha,q,p)(\sigma_{i}\otimes\sigma_{j}))$. The expression for $X_{\chi(\alpha,q,p)}$ is
@@
It quantifies the memory effects, and any non-monotonic increase in $X_{\chi(\alpha,q,p)}$ can be ascribed to quantum information backflow in interval [$q,p$]. The Fig. (6) depicts the variation of quantum memory effects using $X_{\chi(\alpha,q,p)}$. For $q=0.3$ and $\alpha=0.8$, Fig. (6) depicts a non-monotonic increase, implying information backflow which is quantum in nature and not classical [21]. Further non-monotonic variation of $X_{\chi(\alpha,q,p)}$ increases as $\alpha$ increases, indicating enhanced quantum information backflow, hence non-Markovianity. The increase in value of $X_{\chi(\alpha,q,p)}$ is due to the system recovering information, particularly quantum information lost to the environment.
Figure 6: (Color online) The variation of $X_{\chi(\alpha,q,p))}$, Eq. (37), for $q=0.3,\alpha=0.$ (red solid line), $q=0.3,\alpha=0.7$ (blue dashed line), $q=0.3,\alpha=0.8$ (orange dashed-dot line),for $q=0.3,\alpha=0.9$ (purple dashed line), and $q=0.3,\alpha=1$ (green large-dashed line).
@@ IV.1 Non-Markovianity through volume change of physical states of the system
Figure 7: Time evolution of the $|M(p)|$ for a generic Markovian (left graph) and non- Markovian dynamics (right graph). In the Markovian case, the Bloch sphere shrinks along $x$, $y$, and $z$ axis and so the volume decreases monotonically for all values of $p$. For the non-Markovian case, the volume grows suddenly at $p\approx 0.7$ indicating the non-Markovianity beyond this point.
Rate of change in the volume of accessible states to the system is a tool in [22], enabling a geometrical visualization of the dynamical effects of non-Markovianity. Unlike other methods, see, for example, [5], this method need not to be optimized over the initial states. If the quantum evolution is Markovian, then this implies that the domain’s volume of the dynamical map reduces monotonically, as indicated in the left side of Fig. (6). In contrast, for a non-Markovian dynamical map, there could be an increase in the domain’s volume for some time interval during the dynamics.
The density matrix of a qubit can be written using operators $G=\frac{1}{\sqrt{2}}\{I_{0},\sigma_{x},\sigma_{y},\sigma_{z}\}$ and generalized Bloch vector ($\vec{r}$) [3, 25] as
@@
Here, $r_{i}=tr(G_{i}\rho)$ such that $\vec{r}=(\frac{1}{\sqrt{2}},r_{i=1,2,3})$. If a map $\Phi$ acts on a single qubit then $\Phi:\rho\rightarrow\rho^{\prime}$, where $\rho^{\prime}$ must be expressible in terms of a new Bloch vector $\vec{r}^{\prime}$ associated to $\rho^{\prime}$. The action of a unital quantum map $\Phi$ on a qubit state can be represented by mapping the Bloch vector according to $\vec{r}\rightarrow\vec{r}^{\prime}=M\vec{r}$. The equation, $M_{ij}=tr[G_{i}\Phi[G_{j}]]$, gives the respective matrix elements. This is an affine transformation for the Bloch vector, for the respective $M$, as
@@
Here, B is the matrix satisfying condition $|\mathbf{B}|=|M|$. The absolute value of determinant of matrix $M$, $|M|$, can be associated with volume of accessible states [38]. It gives the reduction factor for the volume of accessible states, given by the measure of the set of evolved Bloch vectors, with respect to its value at $p=0$. We plot $|M(p)|$ as a function of $p$ for the evolution of the Bloch vectors under the action of the general depolarizing noisy map in Fig. (7) for $\alpha=0$ (no perturbation), left and $\alpha=0.8$ (with perturbation), right. When $\alpha=0$, the evolution of the norm of the map demonstrates Markovian dynamics, as can be discerned from the monotonic decrease in $|M(p)|$. For $\alpha=0.8$, increase in the value of $|M(p)|$ is observed, for $p$ ranging from $0.7$ to $1.0$,indicating non-Markovianity. In spherical co-ordinates, it is then straightforward to check that any positive trace-preserving map described by Eq. (37) induces the change
@@
where $||\mathbf{B}||$ decreases monotonically for any positive, linear, and trace preserving map [39, 22]. The values of $||\textbf{B}||$ as well as $||M(p)||$ could be used to identify non-Markovianity, as a similar trend like $|M(p)|$ will be observed for them. The $\frac{d||M(p)||}{dp}>0$ for some $p$, is then indicative of non-Markovian dynamics. Using this, the quantification of non-Markovianity for the depolarizing channels is
@@
It is worth mentioning that $\mathcal{N}_{BLP}$ and $\mathcal{N}_{v}$ measures show similarity.

### v2

@@ III.2 The negativity of the decay rate
In some cases of interest, the open system dynamics may be written in terms of a time local master equation involving time-dependent functions as prefactors with otherwise Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) form [60, 61]. Then, for some times during the system evolution, negative decay rates may show up, indicative of the non-Markovianiaty of the dynamics. The generalized Markovian dynamics appears when the master equation takes the quasi-GKSL-form [62]
@@
written in a canonical form [20], where the $L_{i}(p)$ are traceless orthonormal operators. The decay rates $\gamma_{i}(p)\geq 0$ defines a CP divisible dynamical map. It is natural to regard dynamical maps $\Phi(t,0)$ with master equations of the type given by Eq. (III.2) for which $\gamma_{i}(p)<0$ for some $i$ and some time as candidates for non-Markovian quantum dynamics [62]. In these cases, the dynamical map is no longer CP divisible.
@@
The Kraus representation given by Eq. (II) is a solution to the master equation describing the depolarizing map in the canonical form:
@@
Figure 3 illustrates a plot of decay rates $\gamma_{X,Y,Z}(p)$ as a function of $p$ for $\alpha=0.0$ (red dashed curve) and $\alpha=0.7$ (blue solid curve). We observe that when the non-Markovian (perturbation) parameter $\alpha$ is equal to zero, the decay rates become positive for the entire range of $p$ and then the dynamics is CP divisible, as expected. Furthermore, when $\alpha=0.7$, the decay rates turn into negative at the singularity point, $i.e.,\leavevmode\nobreak\ p\approx 0.78$, indicating non-Markovian dynamics.
Using the canonical decay rates $\gamma_{i}(p)$ and the master equation, given by Eq. (18), a function $f_{i}(p)$ is defined [20],
$f_{i}(p)=\max[-\gamma_{i}(p),0]\geq 0.$ (20)
If $f_{i}(p)=0$, at any time, then the evolution is CP divisible. The functions $f_{i}(p)$ can be used to quantify the non-Markovianity as
@@
where $\gamma(p)$ is as in Eq. (19). It diverges because of the singularity of $\gamma(p)$ at $p\approx 0.8$ (see Fig. 3). One solution to this problem, following an idea proposed in [7, 21], is to replace $\gamma(p)$ by its normalized form
@@
=$=\Biggl{\{}\ln\left(2p(-2+\alpha)-4\alpha+3p^{2}\alpha\right)$
$+\frac{6\alpha(\tanh^{-1}\left(\frac{-2+\alpha+3p\alpha}{\sqrt{4-4\alpha+13\alpha^{2}}}\right))}{\sqrt{4-4\alpha+13\alpha^{2}}}\Biggr{\}}^{p=1}_{p=\alpha_{-}}.$ (23)
Figure 4 represents the values of $\mathcal{N}^{{}^{\prime}}_{HCLA}$ (blue solid curve) as a function of non-Markovian (perturbation) parameter $\alpha$. The monotonic increase of this measure with $\alpha$ indicates non-Markovianity. These results are connected to the RHP measure, $\mathcal{N}_{RHP}$, since both coincide for the two level system [20].
Figure 4: A plot of the $BLP$ measure $\mathcal{N}_{BLP}$ (red dashed line) and $HCLA$ measure $\mathcal{N}^{{}^{\prime}}_{HCLA}$ (blue solid curve).
@@ III.3 The distinguishability of quantum states
The BLP method observes the distinguishability between two initial states, $\rho^{S}_{1}$ and $\rho^{S}_{2}$ by calculating the $trace$ $distance$, $D$, between them. $D$ is a metric quantity on the space of physical states, satisfying $0\leq D\leq 1$, where $D=1$ if and only if $\rho^{S}_{1}$ and $\rho^{S}_{2}$ have orthogonal support, and characterizes a realistic upper bound for the distinguishability between the probability distributions resulting from measurements executed on the quantum states. Therefore, the trace distance can be given an interpretation as the distinguishability between two quantum states [5, 3]. If $\Phi(p,0)$ is “CP divisible,” then it reflects as a contraction for the trace distance, $i.e.,$ $D$ is nonincreasing under CP divisible maps,
$D\left[\Phi(p,0)(\rho^{1}_{S}),\Phi(p,0)(\rho^{2}_{S})\right]\leq D\left[\rho^{1}_{S},\rho^{2}_{S}\right],$ (24)
where, in general,
$D\left[\rho^{2}_{S},\rho^{1}_{S}\right]=\frac{1}{2}||\rho^{2}_{S}-\rho^{1}_{S}||_{1}=\frac{1}{2}\text{tr}\sqrt{(\rho^{2}_{S}-\rho^{1}_{S})^{2}}.$ (25)
The violation of Eq. (24) would be an indicator of non-Markovian behavior [5]. This is a sufficient but not necessary condition. The non-Markovian dynamics satisfying this violation is of P-indivisible nature [7]. The violation given by Eq. (24) witnesses non-Markovian effects which are popularly interpreted as information backflow. The information backflow can be quantum as well as classical in nature. We justify, below, the quantumness of information backflow observed in the depolarizing channel, using the recently proposed memory witness in [22].
As an example, we calculate the trace distance between two different initial states, with orthogonal support, which evolve under depolarizing dynamics, given by Eq. (III.1). Let $\rho^{1}_{S}(0)=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix},$ and $\rho^{2}_{S}(0)=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$. By applying the depolarizing channel, given by Eq. (III.1), on these initial states, the final states turn out to be
$\rho^{1}_{S}(t)$ $=\Phi(p,0)\left[\rho_{S}^{1}(0)\right]$
$=\frac{1}{2}\begin{pmatrix}\frac{1}{2}&\frac{p}{2}+\frac{\alpha p}{2}-\frac{3}{8}\alpha p^{2}-\frac{1}{2}\\ \frac{p}{2}+\frac{\alpha p}{2}-\frac{3}{8}\alpha p^{2}-\frac{1}{2}&\frac{1}{2}\end{pmatrix},$
$\rho^{2}_{S}(t)$ $=\Phi(p,0)\left(\rho_{S}^{2}(0)\right)]$
@@
$D\left[\rho^{2}_{S},\rho^{1}_{S}\right]$ $=\frac{1}{2}\text{tr}\sqrt{(\rho^{2}_{S}-\rho^{1}_{S})^{2}}$
@@
Figure 5: Logarithmic plot of trace distance $D$ for $\alpha=0.0$ (solid red curve), $\alpha=0.7$ (blue dash curve), and $\alpha=0.9$ (orange dashed-dotted curve), as a function of $p$. Note that, $\alpha=0.0$ describes the Markovian dynamics since $D$ decreases monotonically for all values of $p$.
The logarithmic plot of trace distance $D$ as a function of $p$ with $\alpha=0$ (solid red curve), $\alpha=0.7$ (blue dashed curve), and $\alpha=0.9$ (orange dash dotted curve), under the considered non-Markovian depolarizing noise is depicted in Fig. 5. Note that larger $\alpha$ shows a larger enhancement region, suggesting larger non-Markovianity in the sense of BLP [5]. The standard form of the BLP measure of non-Markovianity, $\mathcal{N}_{BLP}$, is given by
@@
If $\frac{dD}{dp}>0$ for some $p$, then the dynamics is non-Markovian [7] and the quantification of this non-Markovianity can be done by calculating $\mathcal{N}_{BLP}$. Equation (28) involves maximizing distinguishability across all possible initial states. Recall, the initial states $\rho^{1}_{S}$ and $\rho^{2}_{S}$ in our example possess orthogonal support. Consequently, their distinguishability is at its maximum, which is 1. Hence, the BLP measure is given as
@@
$\mathcal{N}_{BLP}$ is illustrated in Fig. 4 (red dash line) and shows an agreement with the quantification of non-Markovianity according to the normalized $\mathcal{N}^{{}^{\prime}}_{HCLA}$.
@@ III.4 Using quantum memory witness
The intermediate dynamical map $\Phi(p,q)$ and the corresponding Choi matrix $\chi(\alpha,q,p)$ are useful for the observation and certification of non-Markovian memory effects. The breaking of CP divisibility can be captured using the Choi matrix, which is Hermitian, allowing one to write a spectral decomposition as $\chi(\alpha,q,p)=\sum_{i}\lambda_{i}P_{i}$. The $P_{i}$ are orthogonal projections derived using corresponding eigenvectors, and the $\Lambda_{i}$ are eigenvalues associated with $\chi(\alpha,q,p)$. The occurrence of negative $\Lambda_{i}$ points to non-Markovianity. Following this, the Bell states (|$(\ket{\phi^{+}},\ket{\psi^{\pm}})$ are observed to serve as witness operators for the depolarizing channel [64] giving
$\text{tr}(\ket{\phi^{+}}\bra{\phi^{+}}\chi(\alpha,q,p))=\Lambda_{\mathbb{I}},$
$\text{tr}(\ket{\psi^{\pm}}\bra{\psi^{\pm}}\chi(\alpha,q,p))=\Lambda_{i=x,y,z}.$ (30)
One of the eigenvalues ($\Lambda_{\mathbb{I}},\Lambda_{i=x,y,z}$) becomes negative for $q=0.3$ and $\alpha>0.7$, indicating that the intermediate evolution is not CP divisible in the interval [$q,p$] and, hence non-Markovian. The map admits information backflow in this interval and the usual tools lack the ability to distinguish this information as classical or quantum in nature. We use Ref. [22] to characterize the quantum nature of information backflow in the intermediate depolarizing map. The existence of the Hermitian witness operator, in the form of the Bell states, then allows the establishment of the variable $X_{\chi(\alpha,q,p)}$ corresponding to $\chi(\alpha,q,p)$. This is used to ascertain the quantum nature of the information backflow, characterized by nonmonoticity of $X_{\chi(\alpha,q,p)}$, by satisfying the following condition [22]
@@
where $\vec{s}$ consists of $s_{i}=\text{tr}(\chi(\alpha,q,p)(\mathbb{I}\otimes\sigma_{i}))$, and $T$ is a 3$\times$3 matrix consisting of elements $T_{ij}=\text{tr}(\chi(\alpha,q,p)(\sigma_{i}\otimes\sigma_{j}))$. The expression for $X_{\chi(\alpha,q,p)}$ is
@@
It quantifies the memory effects, and any nonmonotonic increase in $X_{\chi(\alpha,q,p)}$ can be ascribed to quantum information backflow in interval [$q,p$]. Fig. 6 depicts the variation of quantum memory effects using $X_{\chi(\alpha,q,p)}$. For $q=0.3$ and $\alpha=0.8$, Fig. 6 depicts a non-monotonic increase, implying information backflow which is quantum in nature and not classical [22]. Further nonmonotonic variation of $X_{\chi(\alpha,q,p)}$ increases as $\alpha$ increases, indicating enhanced quantum information backflow and, hence non-Markovianity. The increase in value of $X_{\chi(\alpha,q,p)}$ is due to the system recovering information, particularly quantum information lost to the environment.
Figure 6: The variation of $X_{\chi(\alpha,q,p))}$, given by Eq. (32), for $q=0.3,\alpha=0.$ (red solid line), $q=0.3,\alpha=0.7$ (blue dashed line), $q=0.3,\alpha=0.8$ (orange dash dotted line), $q=0.3,\alpha=0.9$ (purple dashed line), and $q=0.3,\alpha=1$ (green large-dashed line).
Figure 7: Time evolution of the $|M(p)|$ for a generic Markovian (left graph) and non- Markovian (right graph) dynamics. In the Markovian case, the Bloch sphere shrinks along the $x$, $y$, and $z$ axes and so the volume decreases monotonically for all values of $p$. For the non-Markovian case, the volume suddenly grows at $p\approx 0.7$ indicating the non-Markovianity beyond this point.
@@ IV.1 Non-Markovianity through volume change of physical states of the system
The rate of change in the volume of accessible states to the system is a tool in Ref. [47], enabling a geometrical visualization of the dynamical effects of non-Markovianity. Unlike other methods (see, for example, [5]), this method need not be optimized over the initial states. If the quantum evolution is Markovian, then this implies that the domain’s volume of the dynamical map reduces monotonically, as indicated in the left side of Fig. 7. In contrast, for a non-Markovian dynamical map, there could be an increase in the domain’s volume for some time interval during the dynamics.
The density matrix of a qubit can be written using operators $G=\frac{1}{\sqrt{2}}\{\mathbb{I},\sigma_{X},\sigma_{Y},\sigma_{Z}\}$ and generalized Bloch vector ($\vec{r}$) [3, 65] as
@@
Here, $r_{i}=\text{tr}(G_{i}\rho)$ such that $\vec{r}=(\frac{1}{\sqrt{2}},r_{i=1,2,3})$. If a map $\Phi$ acts on a single qubit then $\Phi:\rho\rightarrow\rho^{\prime}$, where $\rho^{\prime}$ must be expressible in terms of a new Bloch vector $\vec{r}^{\prime}$ associated to $\rho^{\prime}$. The action of a unital quantum map $\Phi$ on a qubit state can be represented by mapping the Bloch vector according to $\vec{r}\rightarrow\vec{r}^{\prime}=M\vec{r}$. The equation, $M_{ij}=\text{tr}[G_{i}\Phi[G_{j}]]$ gives the respective matrix elements. This is an affine transformation for the Bloch vector, for the respective $M$, as
@@
Here, B is the matrix satisfying condition $|\mathbf{B}|=|M|$. The absolute value of the determinant of matrix $M$, $|M|$, can be associated with volume of accessible states [66]. It gives the reduction factor for the volume of accessible states, given by the measure of the set of evolved Bloch vectors, with respect to its value at $p=0$. We plot $|M(p)|$ as a function of $p$ for the evolution of the Bloch vectors under the action of the general depolarizing noisy map in Fig. 7 for $\alpha=0$ (no perturbation; left) and $\alpha=0.8$ (with perturbation), right. When $\alpha=0$, the evolution of the norm of the map demonstrates Markovian dynamics, as can be discerned from the monotonic decrease in $|M(p)|$. For $\alpha=0.8$, increase in the value of $|M(p)|$ is observed, for $p$ ranging from $0.7$ to $1.0$,indicating non-Markovianity. In spherical co-ordinates, it is then straightforward to check that any positive trace-preserving map described by Eq. (32) induces the change
@@
where $||\mathbf{B}||$ decreases monotonically for any positive, linear, and trace preserving map [67, 47]. The values of $||\textbf{B}||$ as well as $||M(p)||_{1}$ could be used to identify non-Markovianity, as a similar trend like $|M(p)|$ will be observed for them. The $\frac{d||M(p)||_{1}}{dp}>0$ for some $p$, is then indicative of non-Markovian dynamics. Using this, the quantification of non-Markovianity for the depolarizing channels is
@@
It is worth mentioning that the $\mathcal{N}_{BLP}$ and $\mathcal{N}_{v}$ measures show similarity.